# SMReferences ℓ

### v1

@@ Introduction
Let $\mathbb{H}$ be the set of Hamiltonians satisfying assumptions (11) and (12). For any $\hat{H}\in\mathbb{H}$, suppose that the following holds.
@@
Proof of the Grabowski–Mathieu conjecture for general boosts
@@
The proof that $\hat{B}$ must take the form of Eq. (14) is given in End Matter. Accepting this fact, we may, without loss of generality, write $\hat{B}=\hat{B}_{\hat{H}}+\hat{R}$.
Since Lemmas 4, 5 (with $\hat{B}_{\hat{H}}$ replaced by $\hat{B}$), and 6 remain valid, it suffices to show $\hat{Q}^{(n)}$ is translationally invariant. If $\hat{Q}^{(k)}$ is translationally invariant, then
@@
where $\hat{X}\coloneqq\mathcal{T}(\hat{R})-\hat{R}$. For $k=2$ this vanishes, proving translational invariance of $\hat{Q}^{(3)}$. For general $k$, the one-locality of $\hat{X}$ again ensures $\hat{X},\hat{Q}^{(k)}]=0$ (see Proposition 12 in End Matter), and hence $\hat{Q}^{(k+1)}$ is also translationally invariant. Thus, all $\hat{Q}^{(n)}$ are translationally invariant, and commutativity follows as in Theorem 2. ∎
@@
The proof is essentially the same as that of Theorem 2, except that commutativity with $\hat{X}$ must also be shown (see End Matter).
Generalization to parametrized Hamiltonians
@@
$\hat{Q}^{(2)}(\lambda)$ $\coloneqq\hat{H}(\lambda),$ (19)
$\hat{Q}^{(n+1)}(\lambda)$ $\coloneqq[\hat{B}_{\hat{H}(\lambda)},\hat{Q}^{(n)}(\lambda)]+(v\cdot\nabla)\hat{Q}^{(n)}(\lambda),$ (20)
@@
$\hat{Q}^{(3)}(\lambda)=-\sum_{i\in\mathbb{Z}}[\hat{h}_{i}(\lambda),\hat{h}_{i+1}(\lambda)]+(v\cdot\nabla)\hat{H}(\lambda),$ (21)
where $\hat{h}_{i}(\lambda)=\hat{h}^{(2)}_{i}(\lambda)+\hat{h}^{(1)}_{i}(\lambda)+\hat{c}^{(1)}_{i}(\lambda)-\hat{c}^{(1)}_{i+1}(\lambda)$ for some translationally invariant $1$-local operator $\hat{C}(\lambda)$. The generalized Reshetikhin condition requires $\hat{Q}^{(3)}(\lambda),\hat{H}(\lambda)]=0$ for all $\lambda\in U$, or its local form [33].

### v3

@@ V Dichotomy on integrability
Let $\mathbb{H}$ be a set of Hamiltonians satisfying assumptions (28) and (29). For any $\hat{H}\in\mathbb{H}$, suppose that the following holds.
@@
VI Generalizations
VI.1 Proof of the Grabowski–Mathieu conjecture for general boosts
@@
The proof that $\hat{B}$ must take the form of Eq. (31) is given in Appendix B.1. Accepting this fact, we may, without loss of generality, write $\hat{B}=\hat{B}_{\hat{H}}+\hat{R}$.
Since Lemmas 7, 8 (with $\hat{B}_{\hat{H}}$ replaced by $\hat{B}$), and 9 remain valid, it suffices to show $\hat{Q}^{(n)}$ is translationally invariant. If $\hat{Q}^{(k)}$ is translationally invariant, then
@@
where $\hat{X}\coloneqq\mathcal{T}(\hat{R})-\hat{R}$. For $k=2$ this vanishes, proving translational invariance of $\hat{Q}^{(3)}$. For general $k$, the one-locality of $\hat{X}$ again ensures $\hat{X},\hat{Q}^{(k)}]=0$ (see Proposition 21 in Appendix B.1), and hence $\hat{Q}^{(k+1)}$ is also translationally invariant. Thus, all $\hat{Q}^{(n)}$ are translationally invariant, and commutativity follows as in Theorem 1. ∎
@@
The proof is essentially the same as that of Theorem 1, except that commutativity with $\hat{X}$ must also be shown (see Appendix B.2).
VI.2 Parametrized Hamiltonians
@@
$\hat{Q}^{(2)}_{v}(\lambda)$ $\coloneqq\hat{H}(\lambda),$ (36)
$\hat{Q}^{(n+1)}_{v}(\lambda)$ $\coloneqq[\hat{B}_{\hat{H}(\lambda)},\hat{Q}^{(n)}_{v}(\lambda)]+(v\cdot\nabla)\hat{Q}^{(n)}_{v}(\lambda),$ (37)
@@
$\hat{Q}^{(3)}_{v}(\lambda)=-\sum_{i\in\mathbb{Z}}[\hat{h}_{i}(\lambda),\hat{h}_{i+1}(\lambda)]+(v\cdot\nabla)\hat{H}(\lambda),$ (38)
where $\hat{h}_{i}(\lambda)=\hat{h}^{(2)}_{i}(\lambda)+\hat{h}^{(1)}_{i}(\lambda)+\hat{c}^{(1)}_{i}(\lambda)-\hat{c}^{(1)}_{i+1}(\lambda)$ for some translationally invariant $1$-local operator $\hat{C}_{v}(\lambda)=\sum_{i\in\mathbb{Z}}\hat{c}^{(1)}_{v,i}(\lambda)$. The generalized Reshetikhin condition requires $\hat{Q}^{(3)}_{v}(\lambda),\hat{H}(\lambda)]=0$ for all $\lambda\in U$, or its local form [38].